OpenStudy (anonymous):

how do i determine the equation of the quadratic function that passes through (-4,5) if its zeroes are 2+√3 and 2-√3

OpenStudy (campbell_st):

start with the perfect square (x -2)^2 = 3 or (x - 2)^2 - 3 = 0 because when the solve for x \[(x -2) = \pm \sqrt{3}\] \[x = 2\pm \sqrt{3}\] then you have the equation y = a((x - 2)^2 - 3) y = a(x^2 - 4x + 1) substitute the point to find a 5 = a((-4)^2 - 4(-4) + 1) 5 = a(16 + 16 + 1) a = 5/33 equation y = 5/33(x^2 - 4x + 1)

OpenStudy (anonymous):

would i do the same if i'm given the x-intercepts √7 and -√7 and the parabola passes through the point (-5,3)

OpenStudy (campbell_st):

something similar that seems easiest to me..... its a solution I worked through, someone may have a more elegant method

OpenStudy (anonymous):

thanks!